# Black Hole Superradiance of Interacting Multi-Field

### v2

@@ Appendix C Stability of quasi-equilibrium
In Sec. IV.1 and Sec. IV.2, we have discussed the asymptotical behavior near $a_{\rm crit}$ and the condition that the quasi-equilibrium becomes unstable. In this appendix, we will show more detailed derivation and generalize our discussion.
@@
However, the quasi-equilibrium stage could be unstable. Considering a small deviation $\delta\epsilon$ from $\epsilon_{\rm eq}$, the evolution of $\delta\epsilon$ is given by
@@
Therefore, we expect that equilibrium breaks down when $\tau_{re}\gtrsim\tau_{var}$, as the evolution of real occupation number $\epsilon$ can not catch up with the change of $\epsilon_{\rm eq}$. In case A, we have $k=1$, and $\tau_{re}<\tau_{var}$ is always satisfied when $a$ is close to ${\tilde{a}}_{\rm crit}$. For $k\geq 1$, there could be threshold in terms of black hole spin ${\tilde{a}}_{th}$, below which one has $\tau_{re}>\tau_{var}$, indicating the equilibrium is unstable. After the black hole spin drops below ${\tilde{a}}_{th}$, the power law decay of $\epsilon_{\rm eq}$ transforms from $t^{-1}$ to $t^{-1/k}$.

### v3

@@ Appendix C Stability of quasi-equilibrium
In Sec. IV.1 and Sec. IV.2, we have discussed the asymptotical behavior near $a_{\rm crit}$ and the condition that the quasiequilibrium becomes unstable. In this appendix, we will show more detailed derivation and generalize our discussion.
@@
However, the quasiequilibrium stage could be unstable. Considering a small deviation $\delta\epsilon$ from $\epsilon_{\rm eq}$, the evolution of $\delta\epsilon$ is given by
@@
Therefore, we expect that equilibrium breaks down when $\tau_{re}\gtrsim\tau_{var}$, as the evolution of real occupation number $\epsilon$ cannot catch up with the change of $\epsilon_{\rm eq}$. In case A, we have $k=1$, and $\tau_{re}<\tau_{var}$ is always satisfied when $a$ is close to ${\tilde{a}}_{\rm crit}$. For $k\geq 1$, there could be threshold in terms of the black hole spin ${\tilde{a}}_{th}$, below which one has $\tau_{re}>\tau_{var}$, indicating the equilibrium is unstable. After the black hole spin drops below ${\tilde{a}}_{th}$, the power-law decay of $\epsilon_{\rm eq}$ transforms from $t^{-1}$ to $t^{-1/k}$.